At what distance (in meters) from a point charge of 7.2×10−16 C does the field produced by this charge have a magnitude of 5.5×10−2 N/C?
Answer #1

here,

the point charge , q = 7.2 * 10^-16 C

let the distance from the charge of that point be r

the magnitude of electric feild , E = K * q /r^2

5.5 * 10^-2 = 9 * 10^9 * 7.2 * 10^-16 /r^2

solving for r

r = 1.09 * 10^-2 m

the distance from the charge of that point is 1.09 * 10^-2 m